# Anomaly detection with flow-based fast calorimeter simulators

### v1

@@ Abstract
Recently, several normalizing flow-based deep generative models have been proposed to accelerate the simulation of calorimeter showers. Using CaloFlow as an example, we show that these models can simultaneously perform unsupervised anomaly detection with no additional training cost. As a demonstration, we consider electromagnetic showers initiated by one (background) or multiple (signal) photons. The CaloFlow model is designed to generate single photon showers, but it also provides access to the shower likelihood. We use this likelihood as an anomaly score and study the showers tagged as being unlikely. As expected, the tagger struggles when the signal photons are nearly collinear, but is otherwise effective. This approach is complementary to a supervised classifier trained on only specific signal models using the same low-level calorimeter inputs. While the supervised classifier is also highly effective at unseen signal models, the unsupervised method is more sensitive in certain regions and thus we expect that the ultimate performance will require a combination of approaches.
@@ I Introduction
Recently, normalizing flows have found successful applications in fast calorimeter simulation tasks [39, 40, 43, 42, 44, 51]. In this paper, we demonstrate the utility of these flow-based generative models as unsupervised anomaly detectors for identifying BSM physics in calorimeter shower data. Specifically, we apply CaloFlow [39], a flow-based fast calorimeter simulation model, to single-photon showers from a new sampling calorimeter version [52] of the CaloGAN dataset [53, 21, 22]. In this context, single-photon showers serve as the background events, while the signal events consist of photon showers originating from a generic BSM particle $\chi$ that undergoes the decay $\chi\to\gamma\gamma$. The $\chi$ particle is taken to be invisible and only interacts indirectly with the calorimeter through its decay products. By training CaloFlow to maximize the log-likelihood when evaluated on background events, we are able to detect the signal events based on a cut on the log-likelihood. We focus on achieving signal sensitivity, but the approach could be combined with a number of background estimation strategies [54, 11].
While we believe this is the first unification of simulation and anomaly detection, both subjects have been well-studied with machine learning. Many deep generative models have been studied for calorimeter simulation [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 35, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50], including a number of proposals developed on the CaloChallenge datasets [55], and they are also now being integrated into experimental workflows [36]. We focus on normalizing flows since they give direct access to the likelihood. This information can also be extracted from diffusion models [56] and it would be interesting to compare approaches in the future. For anomaly detection, unsupervised methods have been extensively studied (e.g. Refs. [57, 58, 59] and many others) and also include density-based approaches [12]. Like the density-based methods, we use the likelihood directly as the anomaly score.111 Note that this is not unique and is sensitive to how the data are represented/preprocessed [60, 61, 62].
@@
This paper is organized as follows. In Sec. II, we describe calorimeter setup and the datasets that were used during training/evaluation. In Sec. III, we explain how CaloFlow is used as an unsupervised anomaly detector by placing cuts on the log-likelihood of background and signal showers evaluated using CaloFlow. In Sec. IV, we include the results of performing unsupervised anomaly detection with CaloFlow. Finally, we conclude in Sec. V.
@@ II Dataset
For this study, we decided to create a new, more realistic sampling calorimeter version [52] of the CaloGAN dataset. The original dataset included energy contributions from both active and inactive calorimeter layers, whereas our new dataset only includes energy contributions from the active layers as would be available in practice. The sampling fraction of our new calorimeter setup is $\sim 20\%$. The simple detector is a three-layer, liquid argon (LAr) sampling calorimeter cube with 480mm side-length that is inspired by the ATLAS LAr electromagnetic calorimeter [63]. It is simulated as follows: Geant4 [64, 65, 66] is used to generate particles and simulate their interaction with our calorimeter using the Ftfp_Bert physics list based on the Fritiof [67, 68, 69, 70] and Bertini intra-nuclear cascade [71, 72, 73] models with the standard electromagnetic physics package [74]. While we use this new simulator to create a dataset for anomaly detection, we expect it should be more generally useful for a broad variety of tasks.
@@ III CaloFlow
CaloFlow [39, 40] is an approach to fast calorimeter simulation based on conditional normalizing flows. In the context of fast calorimeter simulation, CaloFlow uses a two-flow method to learn to generate the voxel level shower energies $\vec{\mathcal{I}}$ conditioned on the corresponding incident energies of the showers $E_{\text{inc}}$ denoted by $p(\vec{\mathcal{I}}|E_{\text{inc}})$. Flow-I is constructed to learn the probability density of calorimeter layer energies [^3] The layer energy of a given calorimeter layer is the sum of all the voxel energies in that layer. $E_{i}$ conditioned on the incident energy $p_{1}(E_{0},E_{1},E_{2}|E_{\text{inc}})$, while Flow-II is designed to learn the probability density of the normalized voxel level shower energies conditioned on the calorimeter layer energies $p_{2}\left(\hat{\mathcal{I}}|E_{0},E_{1},E_{2},E_{\text{inc}}\right)$. By normalized, we mean that the voxel energies in each layer are made to sum to unity. The dimensions of the conditional inputs and outputs of the two flows are shown in Table 2. Importantly, these flows were trained only using single photon showers.
One important difference with original CaloFlow is that in this application to anomaly detection we do not have direct empirical access to $E_{\text{inc}}$. For a given shower, we do not know a priori what the corresponding $E_{\text{inc}}$ is and would instead have to use a reconstructed estimate of $E_{\text{inc}}$ which we denote as $E^{(\text{rec})}_{\text{inc}}$. In this work, we use a simple regression method to reconstruct $E_{\text{inc}}$ given the total deposited energy in the calorimeter $E_{\text{dep}}$. The reconstructed incident energy is defined as $E^{(\text{rec})}_{\text{inc}}=\lambda E_{\text{dep}}$, where $\lambda$ is the mean of $E_{\text{inc}}/E_{\text{dep}}$ in the single photon training dataset. Note that the true $E_{\text{inc}}$ is still used for training CaloFlow, while $E^{(\text{rec})}_{\text{inc}}$ is used when performing anomaly detection.
@@ III.1 Unsupervised anomaly detection with CaloFlow
We see from Figure 2 that the signal showers generally have log-likelihoods that are distinguishable from that of the background showers. Hence, we can use CaloFlow as an unsupervised anomaly detector by placing cuts on the full log-likelihood to discriminate signal from background showers. The results are detailed in Section IV.
@@ IV.1 Decay at fixed displacement
Our performance metric for anomaly detection will be the significance improvement which is defined as the signal efficiency (i.e. true positive rate or tpr) divided by the square root of the background efficiency (i.e. false positive rate or fpr). The maximum significance improvement corresponds to the best possible [^4] This is signal model-dependent, but still provides a useful bound on the achievable performance. multiplicative factor by which the signal significance can be improved with a cut on the anomaly score. Figure 3 shows a heatmap of maximum significance improvement for each of the four different particle masses $m_{\chi}$, and the 10 fixed displacements where the decay occurs. In general, we find that showers from decays at larger $z$ are more anomalous. However, there is clear exception in the case of $m_{\chi}=5$ GeV, where the showers originating from decays that occur before the $\chi$ particle reaches the calorimeter (e.g., $z=$ 0.33 m, 0.66 m) are more anomalous than those from decays occuring at the front face of the calorimeter (i.e., $z=1.0$ m). This is due to the fact that the 5 GeV particle is less boosted compared to the other lighter particles that we consider in this study. As a result, the decay of the 5 GeV particle often results in a wider angle between the produced pair of photons which CaloFlow is better able to distinguish from the background single-photon showers. See Figure 4 for an example of such a shower and comparison with a regular photon shower. On the other hand, if a decay occurs right at the front face of the calorimeter, there is insufficient time for the pair of photons to propagate and create two distinct blobs of energy in the calorimeter.
Figure 4: Top row: Example of shower with two distinct blobs of energy originating from a 5 GeV $\chi$ particle. Bottom row: Example of typical photon shower with centralized energy core. The energy deposition in each of the three layers are shown here with layer 0 on the left, layer 1 in the center, and layer 2 on the right.
@@ IV.2 Varying lifetime
To assess the performance of our unsupervised anomaly detector on realistic scenarios, we construct new datasets consisting of showers from $\chi$ particles with fixed rest frame lifetimes. In particular, for each chosen mass $m_{\chi}$ and lifetime $\tau$, we have a total of 100k events that are made up of showers from $\chi$ decays at fixed displacements. The proportion of showers associated to decays at each fixed displacement is determined based on the lifetime $\tau$. Particles that decay after the calorimeter volume are not included in the events since they are not detected within the calorimeter.
@@
In other words, the number of showers originating from the decay at $z=z_{i}$ is equal to $10^{5}\times\hat{w}_{i}$ [^5] In some cases, we had to round $\hat{w}_{i}$ to ensure that the total number of showers is equal to 100k.. In this work, we consider three possible lifetimes $\tau\in\{0.01,0.1,1\}$ ns for each choice of $m_{\chi}$.
@@
Overall, we find that CaloFlow is able to achieve significantly better performance compared to a random classifier which we take to be the baseline. One exception is for $(m_{\chi},\tau)=(0.5\ \text{GeV},0.01\ \text{ns})$ where the performance closely matches that of the random baseline. The poorer performance here is due to the lower boost factor $\gamma$ for the $0.5$ GeV particle compared to the two lower particles masses which results in the majority of the decays occurring close to the center of the detector ($z=0$ m). Meanwhile, the mass is still low enough that the photons are not widely separated by the time they reach the calorimeter. As we have seen in Figure 3, such showers are mostly not detected as anomalous by CaloFlow. At longer lifetimes, the showers are more anomalous as a larger proportion of them originate from decays occuring within the calorimeter.
For the three lowest $m_{\chi}$, the significance improvement generally increases with background rejection. Also, we find that significance improvement curves are similar across lifetimes for each of these three lowest $m_{\chi}$. At fixed $m_{\chi}$, the large boost of these particles results in a similar proportion of particles decaying at a given fixed displacement before/within the calorimeter for different lifetimes $\tau$.
@@ IV.3 Comparison with supervised anomaly detection
The previous section showed that the unsupervised anomaly detector has broad coverage across the various model parameters. An important question to ask is how this compares to a dedicated search for the $\chi\rightarrow\gamma\gamma$ signal. For a particular signal model, we would expect that model to outperform the unsupervised approach. However, it is not possible to have a dedicated search for every possible signal and so the key question to ask is how well a supervised model trained on one signal would perform on other signals not seen during training.
In this section, we compare the performance of our method against two supervised classifiers. Each supervised classifier was trained on a combined dataset with 100k signal showers and 100k background showers. The showers originate from particles with kinetic energy uniformly distributed in the range [1,100] GeV. The signal showers originate from $\chi$ particles with lifetime $\tau=1.00$ ns. The first (second) supervised classifier was trained on the a dataset with $m_{\chi}=5\times 10^{-3}\ (5)$ GeV. These models were chosen because they are sufficiently different that they would likely be covered by different dedicated searches. It is thus interesting to ask if a search optimized for one of the models would still be sensitive to the other, since the unsupervised approach has some sensitivity to both models.
The supervised classifier is a fully-connected neural network with two hidden layer with 512 nodes each. We have a 505-dimensional input consisting of the voxel energies normalized by the reconstructed incident energy $\vec{\mathcal{I}}/E^{(\text{rec})}_{\text{inc}}$ (504-dim) and the reconstructed incident energy $E^{(\text{rec})}_{\text{inc}}$ (1-dim). The output layer returns a single number which is passed through a sigmoid activation function. All other activation functions are ReLUs. The supervised classifier was trained for a total of 50 epochs with a train/test/validation split of 60%/20%/20%. These parameters were not extensively optimized, but we found little gain from small variations in the setup.
@@
After training the supervised classifiers, they were evaluated on the signal showers discussed in Section IV.2. Figure 6 shows the comparison plots between the performance of our CaloFlow anomaly detector (unsup) and the supervised classifiers (sup). The top (bottom) row shows the performance of the supervised classifier that was trained on signal showers originating from $\chi$ particles with $m_{\chi}=5\times 10^{-3}\ (5)$ GeV and lifetime $\tau=1.00$ ns.
In the top left plot, we see that the supervised classifier clearly performs better than CaloFlow on the signal showers originating from $\chi$ particles that have same mass ($5\times 10^{-3}$ GeV) as what it was trained on, achieving a larger significance improvement for various background rejection. We found that the supervised classifier still performs well on signal showers originating from $\chi$ particles with higher masses as long as the $\chi$ particles they originate from are still highly boosted.
However, as shown in the top right plot, the performance of the supervised classifier decreases significantly for signal showers associated to $m_{\chi}=5$ GeV that are less highly boosted. We find that CaloFlow achieves a much higher significance improvement for large range of background rejection in the cases of $\tau=0.01,0.1$ ns. For the case of $\tau=1$ ns, the supervised classifier and CaloFlow appear to have comparable performance across different background rejection.
In the bottom right plot, we again see that the supervised classifier outperforms CaloFlow on the signal showers originating from $\chi$ particles that have same mass (5 GeV) as what it was trained on. Not surprisingly, we see that this supervised classifier achieves much higher significance improvement compared to the other supervised classifier that was trained on signal showers associated to $m_{\chi}=5\times 10^{-3}$ GeV (top right plot) when evaluated on signal showers associated to $m_{\chi}=5$ GeV.
@@
Similarly, this supervised classifier evaluated on signal showers associated to $m_{\chi}=5\times 10^{-3}$ GeV (bottom left plot) achieves lower significance improvement compared to the supervised classifier that was trained on signal showers associated to $m_{\chi}=5\times 10^{-3}$ GeV (top left plot). Despite the drop in performance, we see that the supervised classifier still outperforms CaloFlow when evaluated on signal showers it was not trained on. This is likely due to CaloFlow not being able to fully discriminate against signal showers that only decay after the first longitudinal layer (i.e., $E_{0}=0$ GeV), whereas such showers are usually perfectly distinguished by the supervised classifier. As noise is added to the voxel energies when using CaloFlow, this artificially causes $E_{0}>0$ GeV [^6] We checked that a different treatment of the layers with zero energy deposition does not improve the performance. and prevents such showers from being perfectly distinguished by CaloFlow.

### v2

@@ Abstract
Recently, several normalizing flow-based deep generative models have been proposed to accelerate the simulation of calorimeter showers. Using CaloFlow as an example, we show that these models can simultaneously perform unsupervised anomaly detection with no additional training cost. As a demonstration, we consider electromagnetic showers initiated by one (background) or multiple (signal) photons. The CaloFlow model is designed to generate single-photon showers, but it also provides access to the shower likelihood. We use this likelihood as an anomaly score and study the showers tagged as being unlikely. As expected, the tagger struggles when the signal photons are nearly collinear, but is otherwise effective. This approach is complementary to a supervised classifier trained on only specific signal models using the same low-level calorimeter inputs. While the supervised classifier is also highly effective at unseen signal models, the unsupervised method is more sensitive in certain regions and thus we expect that the ultimate performance will require a combination of these approaches.
@@ I Introduction
Recently, normalizing flows have found successful applications in fast calorimeter simulation tasks [21, 22, 23, 24, 25, 26, 27]. Moreover, normalizing flows have demonstrated comparably excellent performance across various tasks within high energy physics [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 42, 47, 48, 49, 50, 51, 52, 53]. In this paper, we demonstrate the utility of these flow-based generative models as unsupervised anomaly detectors for identifying BSM physics in calorimeter shower data. Specifically, we apply CaloFlow [21], a flow-based fast calorimeter simulation model, to single-photon showers from a new sampling calorimeter version [54] of the CaloGAN dataset [55, 56, 57]. In this context, single-photon showers serve as the background events, while the signal events consist of photon showers originating from a generic BSM particle $\chi$ that undergoes the decay $\chi\to\gamma\gamma$. The $\chi$ particle is taken to be invisible and interacts only indirectly with the calorimeter through its decay products. By training CaloFlow to maximize the log-likelihood when evaluated on background events, we are able to detect the signal events based on a cut on the log-likelihood. We focus on achieving signal sensitivity, but the approach could be combined with a number of background estimation strategies [28, 11].
While we believe this is the first unification of simulation and anomaly detection, both subjects have been well-studied with machine learning. Many deep generative models have been studied for calorimeter simulation [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 70, 72, 73, 21, 22, 74, 24, 23, 25, 75, 76, 77, 78, 79, 80, 81, 27], including a number of proposals developed on the CaloChallenge datasets [82], and they are also now being integrated into experimental workflows [71]. We focus on normalizing flows since they give direct access to the likelihood. This information can also be extracted from diffusion models [83] and it would be interesting to compare approaches in the future. For anomaly detection, unsupervised methods have been extensively studied (e.g. Refs. [84, 85, 86] and many others) and also include density-based approaches [12]. Like the density-based methods, we use the likelihood directly as the anomaly score.111 Note that this is not unique and is sensitive to how the data are represented/preprocessed [87, 88, 89].
@@
This paper is organized as follows. In Sec. II, we describe the calorimeter setup and the datasets that were used during training/evaluation. In Sec. III, we explain how CaloFlow is used as an unsupervised anomaly detector by placing cuts on the log-likelihood of background and signal showers evaluated using CaloFlow. In Sec. IV, we include the results of performing unsupervised anomaly detection with CaloFlow. Finally, we conclude in Sec. V.
@@ II Dataset
For this study, we decided to create a new, more realistic sampling calorimeter version [54] of the CaloGAN dataset. The original dataset included energy contributions from both active and inactive calorimeter layers, whereas our new dataset includes only energy contributions from the active layers as would be available in practice. The sampling fraction of our new calorimeter setup is $\sim 20\%$. The simple detector is a three-layer, liquid argon (LAr) sampling calorimeter cube with 480mm side length that is inspired by the ATLAS LAr electromagnetic calorimeter [90]. It is simulated as follows: Geant4 [91, 92, 93] is used to generate particles and simulate their interaction with our calorimeter using the Ftfp_Bert physics list based on the Fritiof [94, 95, 96, 97] and Bertini intranuclear cascade [98, 99, 100] models with the standard electromagnetic physics package [101]. While we use this new simulator to create a dataset for anomaly detection, we expect it should be more generally useful for a broad variety of tasks.
@@ III CaloFlow
CaloFlow [21, 22] is an approach to fast calorimeter simulation based on conditional normalizing flows. In the context of fast calorimeter simulation, CaloFlow uses a two-flow method to learn to generate the voxel level shower energies $\vec{\mathcal{I}}$ conditioned on the corresponding incident energies of the showers $E_{\text{inc}}$ denoted by $p(\vec{\mathcal{I}}|E_{\text{inc}})$. Flow-I is constructed to learn the probability density of calorimeter layer energies [^3] The layer energy of a given calorimeter layer is the sum of all the voxel energies in that layer. $E_{i}$ conditioned on the incident energy $p_{1}(E_{0},E_{1},E_{2}|E_{\text{inc}})$, while Flow-II is designed to learn the probability density of the normalized voxel level shower energies conditioned on the calorimeter layer energies and incident energies $p_{2}\left(\hat{\mathcal{I}}|E_{0},E_{1},E_{2},E_{\text{inc}}\right)$. By normalized, we mean that the voxel energies in each layer are made to sum to unity. The dimensions of the conditional inputs and outputs of the two flows are shown in Table 2. Importantly, these flows were trained only using single photon showers.
One important difference from the original CaloFlow is that in this application to anomaly detection we do not have direct empirical access to $E_{\text{inc}}$. For a given shower, we do not know a priori what the corresponding $E_{\text{inc}}$ is and would instead have to use a reconstructed estimate of $E_{\text{inc}}$ which we denote as $E^{(\text{rec})}_{\text{inc}}$. In this work, we use a simple regression method to reconstruct $E_{\text{inc}}$ given the total deposited energy in the calorimeter $E_{\text{dep}}$. The reconstructed incident energy is defined as $E^{(\text{rec})}_{\text{inc}}=\lambda E_{\text{dep}}$, where $\lambda$ is the mean of $E_{\text{inc}}/E_{\text{dep}}$ in the single photon training dataset. Note that the true $E_{\text{inc}}$ is still used for training CaloFlow, while $E^{(\text{rec})}_{\text{inc}}$ is used when performing anomaly detection.
@@ III.1 Unsupervised anomaly detection with CaloFlow
We see from Figure 2 that the signal showers generally have log-likelihoods that are distinguishable from that of the background showers. Hence, we can use CaloFlow as an unsupervised anomaly detector by placing cuts on the full log-likelihood to discriminate signal from background showers. The results are detailed in Section IV. Though not explained here, the main features found in Figure 2 can be understood from the discussion in Section IV.1.
Despite the possibility of data-MC differences affecting the sensitivity of our approach, we note that the accuracy would not be affected since one would presumably apply standard downstream background estimation. Furthermore, experiments usually calibrate their fast simulation and those calibrations could be applied to improve the sensitivity.
@@ IV.1 Decay at fixed displacement
Our performance metric for anomaly detection will be the significance improvement which is defined as the signal efficiency (i.e. true positive rate or tpr) divided by the square root of the background efficiency (i.e. false positive rate or fpr). The maximum significance improvement corresponds to the best possible [^4] This is signal model dependent, but still provides a useful bound on the achievable performance. multiplicative factor by which the signal significance can be improved with a cut on the anomaly score. Figure 3 shows a heatmap of maximum significance improvement for each of the four different particle masses $m_{\chi}$, and the 10 fixed displacements where the decay occurs. In general, we find that showers from decays at larger $z$ are more anomalous. However, there is a clear exception in the case of $m_{\chi}=5$ GeV, where the showers originating from decays that occur before the $\chi$ particle reaches the calorimeter (e.g., $z=$ 0.33 m, 0.66 m) are more anomalous than those from decays occurring at the front face of the calorimeter (i.e., $z=1.0$ m). This is due to the fact that the 5 GeV particle is less boosted compared to the other lighter particles that we consider in this study. As a result, the decay of the 5 GeV particle often results in a wider angle between the produced pair of photons which CaloFlow is better able to distinguish from the background single-photon showers. See Figure 4 for an example of such a shower and comparison with a regular photon shower. On the other hand, if a decay occurs right at the front face of the calorimeter, there is insufficient time for the pair of photons to propagate and create two distinct blobs of energy in the calorimeter. The sudden jump in maximum significance improvement for all $m_{\chi}$ when going from $z=1.32$ m to $z=1.4$ m is due to the discretization of the calorimeter voxel geometry in the longitudinal direction described in Section II. The transition between the second and third longitudinal layers occurs at $z=1.437$ m. Hence, showers from decays at $z=1.4$ m, which have more energy deposited in the third layer, are flagged as significantly more anomalous.
Figure 4: Top row: Example of a shower with two distinct blobs of energy originating from a 5 GeV $\chi$ particle that decayed at $z<1.0$ m. Bottom row: Example of a typical photon shower with centralized energy core. The energy deposition in each of the three layers is shown here with layer 0 on the left, layer 1 in the center, and layer 2 on the right.
@@ IV.2 Varying lifetime
To assess the performance of our unsupervised anomaly detector on realistic scenarios, as mentioned in Section II, we construct new datasets consisting of showers from $\chi$ particles with fixed rest frame lifetimes. In particular, for each chosen mass $m_{\chi}$ and lifetime $\tau$, we have a total of 100k events that are made up of showers from $\chi$ decays at fixed displacements. The proportion of showers associated to decays at each fixed displacement is determined based on the lifetime $\tau$. Particles that decay after the calorimeter volume are not included in the events since they are not detected within the calorimeter.
@@
In other words, the number of showers originating from the decay at $z=z_{i}$ is equal to $10^{5}\times\hat{w}_{i}$ [^5] In some cases, we had to round $\hat{w}_{i}$ to ensure that the total number of showers is equal to 100k.. In this work, we consider three possible lifetimes $\tau\in\{0.01,0.1,1\}$ ns for each choice of $m_{\chi}$. A visualization of how the fixed lifetime datasets are constructed from showers in the fixed displacement datasets is shown in Figure 5. In Table 3, we show the percentage of $\chi$ decays that occur before/within the calorimeter volume and also the average flight length of the $\chi$ particle for different $m_{\chi}$ and lifetimes $\tau$.
Figure 5: Visualization of how the fixed lifetime datasets are constructed by sampling from the fixed displacement datasets. The fraction of total showers in each fixed lifetime dataset originating from $\chi$ decay at displacement $z_{i}$ is set to be $\hat{w}_{i}$.
$\tau$ $m_{\chi}$ $5\times 10^{-3}$ GeV $5\times 10^{-2}$ GeV $5\times 10^{-1}$ GeV $5$ GeV
0.01 ns 5% / 30 m 38% / 3 m 99% / 0.3 m 100% / 0.03 m
0.1 ns 0.5% / 300 m 5% / 30 m 38% / 3 m 99% / 0.3 m
1 ns 0.05% / 3000 m 0.5% / 300 m 5% / 30 m 38% / 3 m
Table 3: Percentage of $\chi$ decays (left number) that occur before/within the calorimeter volume, and the average flight length (right number) of $\chi$ particles for different $m_{\chi}$ and lifetimes $\tau$.
@@
Overall, we find that CaloFlow is able to achieve significantly better performance compared to a random classifier which we take to be the baseline. One exception is for $(m_{\chi},\tau)=(0.5\ \text{GeV},0.01\ \text{ns})$ where the performance closely matches that of the random baseline. The poorer performance here is due to the lower boost factor $\gamma$ for the $0.5$ GeV particle compared to the two lower particles masses which results in the majority of the decays occurring close to the center of the detector ($z=0$ m). Meanwhile, the mass is still low enough that the photons are not widely separated by the time they reach the calorimeter. As we have seen in Figure 3, such showers are mostly not detected as anomalous by CaloFlow. At longer lifetimes, the showers are more anomalous as a larger proportion of them originate from decays occurring within the calorimeter.
For the three lowest $m_{\chi}$, the significance improvement generally increases with background rejection. Also, we find that the significance improvement curves are similar across lifetimes for the two lowest $m_{\chi}$. In these cases, $\frac{z}{c\tau\sqrt{\gamma^{2}-1}}$ is small which implies that $P_{d}(z)\approx\frac{z}{c\tau\sqrt{\gamma^{2}-1}}$. Hence, the lifetime $\tau$ cancels out when computing $\hat{w}_{i}$. In other words, at fixed $m_{\chi}$, the large boost of these particles results in a similar proportion of particles decaying at a given fixed displacement before/within the calorimeter for different lifetimes $\tau$.
@@
There is a local maximum in each of the significance improvement curves for $m_{\chi}=5$ GeV. To understand the local maximum, we have to look at the log-likelihood plot for $m_{\chi}=5$ GeV shown in Figure 7. The peaks at the log-likelihoods of -1000 and -500 are due to showers from decays occurring at $z=$ 0.33 m and 0.66 m respectively which can also be seen in Figure 2. Sliding the cut on log-likelihood in the direction of decreasing log-likelihood is equivalent to increasing the background rejection (1/fpr). Notice that sliding the cut in the direction of decreasing log-likelihood across peaks in the $\chi$ shower log-likelihood curve would create a local maximum in the plot of significance improvement vs background rejection (1/fpr) since the tpr decreases faster than the increase in $1/\sqrt{\rm fpr}$.
Figure 7: Plot of log-likelihood of $\chi$ showers for $m_{\chi}=5$ GeV at the different $\chi$ particle lifetimes.
@@ IV.3 Comparison with supervised anomaly detection
The previous section showed that the unsupervised anomaly detector has broad coverage across the various model parameters. An important question to ask is how this compares to a dedicated search for the $\chi\rightarrow\gamma\gamma$ signal. For a particular signal model, we would expect the dedicated search to outperform the unsupervised approach. However, it is not possible to have a dedicated search for every possible signal and so the key question to ask is how well a supervised model trained on one signal would perform on other signals not seen during training.
In this section, we compare the performance of our method against two supervised classifiers. Each supervised classifier was trained on a combined dataset with 100k signal showers and 100k background showers. The showers originate from particles with kinetic energy uniformly distributed in the range [1,100] GeV. The signal showers originate from $\chi$ particles with lifetime $\tau=1.00$ ns. The first (second) supervised classifier was trained on a dataset with $m_{\chi}=5\times 10^{-3}\ (5)$ GeV. These models were chosen because they are sufficiently different that they would likely be covered by different dedicated searches. It is thus interesting to ask if a search optimized for one of the models would still be sensitive to the other, since the unsupervised approach has some sensitivity to both models.
The supervised classifier is a fully-connected neural network with two hidden layer with 512 nodes each. We have a 505-dimensional input consisting of the voxel energies normalized by the reconstructed incident energy $\vec{\mathcal{I}}/E^{(\text{rec})}_{\text{inc}}$ (504-dim) and the reconstructed incident energy $E^{(\text{rec})}_{\text{inc}}$ (1-dim). The output layer returns a single number which is passed through a sigmoid activation function. All other activation functions are ReLUs. The supervised classifier was trained for a total of 50 epochs with a train/test/validation split of 60%/20%/20%. These parameters were not extensively optimized, but we found little gain from small variations in the setup. CaloFlow has a total of $\sim 38$M parameters, while the supervised classifier only has $\sim 522$k parameters. However, we found that increasing the number of parameters for the supervised classifier does not help to increase its performance. CaloFlow has a larger number of parameters because it has a more difficult task of learning the likelihood of the data.
Top row: Supervised classifier trained on $\mathbf{m_{\chi}=5\times 10^{-3}}$ GeV
Refer to captionRefer to caption
Bottom row: Supervised classifier trained on $\mathbf{m_{\chi}=5}$ GeV
Refer to captionRefer to caption
@@
After training the supervised classifiers, they were evaluated on the signal showers discussed in Section IV.2. Figure 8 shows the comparison plots between the performance of our CaloFlow anomaly detector (unsup) and the supervised classifiers (sup). We see from the figure that the supervised classifier always outperforms CaloFlow on showers originating from $\chi$ particles that have the same mass as what it was trained on (see top left and bottom right plots). However, the performance of the supervised classifier usually decreases when evaluated on signal showers originating from $\chi$ particles that have very different masses from what it was trained on (see top right and bottom left plots). Whether the supervised classifier or CaloFlow achieves better performance depends on $m_{\chi}$.
  • For the upper right plot (trained on $m_{\chi}=5\times 10^{-3}$ GeV and evaluated on $m_{\chi}=5$ GeV), CaloFlow easily outperforms the supervised classifier as the supervised classifier did not see training examples of anomalous showers that are characteristic of larger $m_{\chi}$ (e.g. early decay resulting in two blobs) and is unable to generalize its performance.
@@
  • However, in the lower left plot (trained on $m_{\chi}=5$ GeV and evaluated on $m_{\chi}=5\times 10^{-3}$ GeV) CaloFlow does not outperform the supervised classifier. This is likely due to CaloFlow not being able to fully discriminate against signal showers that only decay after the first longitudinal layer (i.e., $E_{0}=0$ GeV), whereas such showers are usually perfectly distinguished by the supervised classifier. [^6] Here, the point is that CaloFlow is not able to distinguish showers from late decays as well as the supervised classifier can. Nevertheless, for $m_{\chi}=5\times 10^{-3}$ GeV, CaloFlow still has a better anomaly detection performance on late decay (larger $z$) showers compared to early decay (smaller $z$) showers. As noise is added to the voxel energies when using CaloFlow, this artificially causes $E_{0}>0$ GeV [^7] We checked that a different treatment of the layers with zero energy deposition does not improve the performance. and prevents such showers from being perfectly distinguished by CaloFlow.
Finally, let us also comment on some other interesting features in Fig. 8. When the classifier or anomaly detector is evaluated on $m_{\chi}=5\times 10^{-3}$ GeV (left column in Fig. 8), the performance is the same for all the lifetimes considered. This is because, for $m_{\chi}=5\times 10^{-3}$ GeV, the particles are highly boosted, so there is a similar proportion of the total number particles decaying at a given fixed displacement before/within the calorimeter for different lifetimes. (Keep in mind that we only consider particles decaying before/within the calorimeter.) Interestingly, the fully supervised classifier is also the same for all lifetimes for the bottom right plot (trained and evaluated on $m_{\chi}=5$ GeV). Here, the reason is that the signal showers look extremely different from the background photon showers and are perfectly distinguished by the classifier. Thus, we see that the significance improvement is equal to $1/\sqrt{\rm fpr}$.